5. A 130 N force is applied to a 7 kg block. The 7 kg block pushes on a 3 kg block. Both of the blocks accelerate to the right as shown below (the 3 kg block does not fall down). There is nothing connecting the 3 kg block to the 7 kg block. If the 3 kg block is on the verge of slipping and falling, what is the coefficient of friction between the 7 kg block and the 3 kg block?
(a) 0.33
(b) 0.49
(c) 0.61
(d) 0.75
1
Expert's answer
2020-11-10T06:57:30-0500

"F=(m+M)a\\a=\frac{F}{m+M}\\ma=\mu mg\\\mu g=a=\frac{F}{m+M}\\9.8\mu=\frac{130}{3+7}\\\mu=1.33"